OpenStudy (anonymous):

Given f(x)=9x^9, find f^-1(x). Then state whether f^-1(x) is a function.

OpenStudy (anonymous):

\[\Large f(x)=9x^9 \Leftrightarrow y=9x^9 \] to find the inverse, swap the x and the y and solve again for the given form: \[\Large x=9y^9 \] Can you solve for y?

OpenStudy (anonymous):

I get to y^9=x/9 and I'm not sure what id do from there

OpenStudy (anonymous):

take the 9th root :)

OpenStudy (anonymous):

idk what the 9th root of x/9 would be

OpenStudy (anonymous):

it's a value, written as: \[\Large y=\sqrt[9]{\frac{x}{9}} \]

OpenStudy (anonymous):

but that's not an answer choice

OpenStudy (anonymous):

you can list me down your answer choices if you want, however, that is the inverse function .

OpenStudy (anonymous):

these are my choices

OpenStudy (anonymous):

\[\frac{ x }{ 9 }^{1/9}\] is the answer

OpenStudy (anonymous):

and f^-1(x) is a function so c is your answer

OpenStudy (anonymous):

thank you